# ChatBench: From Static Benchmarks to Human-AI Evaluation

### v1

@@ 1 Introduction
In 2024, nearly 40% of US adults reported using generative AI in their everyday lives, an unprecedented rate of adoption for a new technology Bick et al. (2024). As these models, particularly large language models (LLMs), become more integrated into our lives, it becomes increasingly important to evaluate them based on not only their capabilities in isolation, but also their interactions with humans. However, there is a large gap between human interactions and how standard benchmarks, such as Massive Multitask Language Understanding (MMLU), evaluate models Hendrycks et al. (2021). These benchmarks test models on a fixed set of questions, and for each question, they prompt the model with the entire question text and often constrain it to respond with a single multiple choice option as its answer. In contrast, interactions with human users are far more variable, open-ended, and subject to ambiguity. Even conditioned on the same underlying intent, users may phrase their prompts differently, leave out information in their early prompts, or rely on context in later prompts. Robust AI models need to understand how to work with users in these contexts to provide accurate information and complement human expertise.
@@ 3.1 Task Flow and Interface
Figure 1 shows the flow of our user study. In Phase 1, users are asked to answer each question to the best of their ability on their own. In Phase 2, users are asked to chat with an unnamed “AI Chatbot” to help them answer their question. We test two LLMs, contrasting GPT-4o as a strong model and Llama-3.1-8b as a relatively weaker model. We require interaction in Phase 2—the user cannot move onto the next question without sending a message and we say that low-effort conversations, e.g., only “hi”, will be flagged—but otherwise, we do not specify at all how the user should interact with the AI Chatbot. In both phases, users are asked to first report how confident they are about approaching the problem, before attempting to answer it. This additional question-level variable allows us to analyze how AI assistance helps users across varying levels of confidence. After Phase 2, all users provide feedback on the task, with free-text responses including whether they found the AI Chatbot helpful and if they saw it make any mistakes. In Figure 2, we show a screenshot of what users see in Phase 2; in the Appendix, we provide screenshots of all other pages in our task (Figures A2-A9).
@@ 4.3 Characterizing User-AI Conversations
Our summary results show that user-AI accuracies are significantly different from AI-alone accuracies. To better understand what drives these differences we use a separate LLM as an annotator to characterize the user-AI conversations. For each user-AI conversation, we gather the full log of the conversation and its associated metadata (e.g., the question ID, the correct answer, the user’s selected answer, etc.), and prompt a separate instance of GPT-4o to use this information to extract the answers to several classification questions: whether the first substantive user prompt was a question, whether the first user question was a near-exact rephrasing of the original question or one of several other possibilities, and whether the first and last AI answers were correct (Listing 5).
@@
How often does the conversation follow what we might expect if AI benchmarks were faithful proxies of human-AI interaction? We say a conversation mirrors an AI benchmark if (1) the user’s first substantive prompt is a near-exact rephrasing of the benchmark question (otherwise the user is injecting their own knowledge or information into the interaction), (2) the LLM responds with an answer, and (3) the user submits that answer. In Figure 5, we see that only $34\%$ of all interactions mirror AI benchmarks, revealing the extent to which user-AI interactions diverge from AI benchmarks.
@@
First, we study cases where user interaction introduces errors, by focusing on questions where AI-alone free-text is always correct (accuracy of 100% over 50 trials) but the user-AI interaction results in the wrong answer. This could happen either if the user decided to ignore the AI’s correct answer (e.g., if they believed they knew the answer or due to lack of effort) or if the change from AI-alone prompting to user prompting resulted in the AI no longer providing the correct answer. We find much more evidence for the latter. Among over 300 of these interactions with GPT-4o, the model only provided the correct answer in 36% of interactions, and in the remaining interactions, the model either did not provide any clear answer (43%) or provided a wrong answer (21%). Among 116 interactions with Llama-3.1-8b (there are fewer, since there are fewer questions where Llama-3.1-8b achieves perfect accuracy on its own), the model only provided the correct answer in 23% of interactions, instead providing the wrong answer in 43% and no clear answer in 34% of interactions.
We also find that in the majority of these interactions (69% for GPT-4o and 58% for Llama-3.1-8b), the user’s first substantive prompt is not a near-exact rephrasing of the benchmark question, providing further evidence for our hypothesis that the change in accuracy is largely due to the shift in prompting from AI-alone benchmarks to human users. We find that a primary source of divergence is the user asking a related but different question, which is often ambiguous (e.g., leaving out critical information for a math problem).
@@
Next, we study the opposite scenario: questions where AI-alone free-text’s accuracy is poor (below 10% over 50 trials) but the user-AI interaction arrives at the correct answer. We find that in nearly half of these interactions, the AI model actually provided the correct answer to the user (49% for GPT-4o and 43% for Llama-3.1-8b), suggesting that the user’s prompting enabled the AI to arrive at the correct answer, even though it could not on its own. Notably, we find that in around 10% of these interactions, over the course of the interaction the AI improves, from providing either an incorrect answer or no clear answer to providing the correct answer, demonstrating the possibility for AI to improve through user interaction and the need for multi-turn analysis beyond static benchmarks.
Even when the AI is not able to arrive at the correct answer, we find that users are sometimes still able to correct the mistake and select the right answer. We visualize the rates of these corrections in Figure B1, showing that this occurs especially often with the weaker model, Llama-3.1-8b.
@@ 5.1 Fine-Tuning a User Simulator
We then transform the real user-AI conversations from our study into training examples for supervised fine-tuning. Each conversation with $k$ user utterances yields $k+1$ training examples: one example in the Task 1 format where the gold standard response is the real user’s first utterance; $k-1$ examples in the Task 2 format where the gold standard response is each of the remaining utterances (providing the conversation up to that utterance); and one example in the Task 2 format with the full conversation and the gold standard response being “Answer: LETTER” corresponding to the user’s selected multiple choice option.
@@ Results.
As shown in Table 1, fine-tuning our simulator yields large gains, with a 22-26 point increase in correlation and a 21-26% decrease in MAE. As shown in Figures B3-B4, a primary failure mode of the simulator before fine-tuning is that it cannot replicate human mistakes and greatly overestimates user-AI performance, producing far more questions with accuracies of 1.0 than we see in the real user-AI distribution, while the fine-tuned simulator matches the real distribution more closely. We also find that fine-tuning improves the realism of the simulator’s generated utterances, with 11-16 point improvements in BLEU and ROUGE. The fine-tuned simulator also outperforms both AI-alone methods and IQA-EVAL across metrics.
@@ 8 Broader Impacts and Ethical Considerations
Our work is driven by broader impacts: we seek to make AI evaluation more realistic and human-centered, by investigating how evaluation conclusions change when we incorporate human interactions. With our carefully designed user study, we show that evaluation conclusions change significantly from AI-alone to user-AI settings (for the same set of questions), and these results hold over different subject areas, AI models, AI-alone methods, and user-AI conditions. We hope that our work motivates AI researchers and practitioners to think more carefully about human-AI interactions when they evaluate AI systems, instead of only using AI-alone benchmarks.
@@ Acknowledgments
The authors thank Rich Ciapala for computing support and Emma Pierson for helpful discussions.
@@ Appendix A Details on User Study
We recruited workers on Prolific to participate in our study. All Prolific workers who were located in the US, fluent in English, and had not participated in one of our pilots were eligible for our study. We used Prolific’s standard sample, which distributed our study to available participants. Based on early pilots, we estimated that the task took around 25 minutes. We paid all participants \$5.00 upon completion of the entire task. We experimented with offering a small bonus per correct answer, which we discuss below.
A.1 Pilots and Incentivization

### v2

@@ 1 Introduction
In 2024, nearly 40% of US adults reported using generative AI in their everyday lives, an unprecedented rate of adoption for a new technology Bick et al. (2024). As these models, particularly large language models (LLMs), become more integrated into our lives, it becomes increasingly important to evaluate not only their capabilities in isolation, but also on how effectively they support people in solving a variety of tasks. However, there is a large gap between human interactions and how standard benchmarks, such as Massive Multitask Language Understanding (MMLU), evaluate models Hendrycks et al. (2021). These benchmarks test models on a fixed set of questions, and for each question, they prompt the model with the entire question text and often constrain it to respond with a single multiple choice option as its answer. In contrast, interactions with human users are far more variable, open-ended, and subject to ambiguity. Even conditioned on the same underlying intent, users may phrase their prompts differently, leave out information in their early prompts, or rely on context in later prompts. Robust AI models must be designed to handle user interactions in these contexts to provide accurate information and complement human expertise.
@@ 3.1 Task Flow and Interface
Figure 1 shows the flow of our user study. In Phase 1, users are asked to answer each question to the best of their ability on their own. In Phase 2, users are asked to chat with an unnamed “AI Chatbot” to help them answer their question. We test two LLMs, contrasting GPT-4o as a strong model and Llama-3.1-8b as a relatively weaker model. We require interaction in Phase 2—the user cannot move onto the next question without sending a message and we say that low-effort conversations, e.g., only “hi”, will be flagged—but otherwise, we do not specify at all how the user should interact with the AI Chatbot. In both phases, users are asked to first report how confident they are about approaching the problem, before attempting to answer it. This additional question-level variable allows us to analyze how AI assistance helps users across varying levels of confidence. After Phase 2, all users provide feedback on the task, with free-text responses including whether they found the AI Chatbot helpful and if they saw it make any mistakes. In Figure 2, we show a screenshot of what users see in Phase 2; in the Appendix, we provide screenshots of all other pages in our task (Figures A3-A10).
@@ 4.3 Characterizing User-AI Conversations
Our summary results show that user-AI accuracies are significantly different from AI-alone accuracies. To better understand what drives these differences we use a separate LLM as an annotator to characterize the user-AI conversations. For each user-AI conversation, we gather the full log of the conversation and the correct answer to the question, and prompt a separate instance of GPT-4o to use this information to extract the answers to several classification questions: 1) whether the first user question was a near-exact rephrasing of the original question to which an intelligent person or AI would respond with the correct answer among the answer choices; 2) whether the first AI answer is correct; 3) whether the AI provides an answer to the question at hand more than once; and 4) which of the four answer choices the AI’s first answer corresponds to, or “none” if there is no such answer (Listing 5).
@@
How often does the conversation follow what we might expect if AI benchmarks were faithful proxies of human-AI interaction? We say a conversation mirrors an AI benchmark if (1) the user’s first substantive prompt is a near-exact rephrasing of the benchmark question (otherwise the user is injecting their own knowledge or information into the interaction), (2) the LLM responds with only a single answer during the entire interaction, and (3) the user submits that answer. In Figure 5, we see that only $39.8\%$ of all interactions mirror AI benchmarks, revealing the extent to which user-AI interactions diverge from AI benchmarks.
@@
First, we study cases where user interaction introduces errors, by focusing on questions where AI-alone free-text is always correct (accuracy of 100% over 50 trials) but the user-AI interaction results in the wrong answer. This could happen either if the user decided to ignore the AI’s correct answer (e.g., if they believed they knew the answer or due to lack of effort) or if the change from AI-alone prompting to user prompting resulted in the AI no longer providing the correct answer. We find much more evidence for the latter. Among over 300 of these interactions with GPT-4o, the model only provided the correct answer in 26% of interactions, and in the remaining interactions, the model either did not provide any clear answer (67%) or provided a wrong answer (7%). Among 122 interactions with Llama-3.1-8b (there are fewer because there are fewer questions where Llama-3.1-8b achieves perfect accuracy on its own), the model only provided the correct answer in 20% of interactions, instead providing no clear answer in 68% of interactions and the wrong answer in 12% of interactions.
We also find that in the majority of these interactions (66.2% for GPT-4o and 59.8% for Llama-3.1-8b), the user’s first substantive prompt is not a near-exact rephrasing of the benchmark question, providing further evidence for our hypothesis that the change in accuracy is largely due to the shift in prompting from AI-alone benchmarks to human users. We find that a primary source of divergence is the user asking a related but different question, which is often ambiguous (e.g., leaving out critical information for a math problem).
@@
Next, we study the opposite scenario: questions where AI-alone free-text’s accuracy is poor (below 10% over 50 trials) but the user-AI interaction arrives at the correct answer (there are approximately 100 such interactions for each model). In a substantial fraction of these interactions, the AI model actually provided the correct answer to the user (40% for GPT-4o and 32% for Llama-3.1-8b), suggesting that the user’s prompting enabled the AI to arrive at the correct answer, even though it could not on its own. Notably, we find that in around 10% of these interactions the AI’s responses shift from incorrect or unclear to correct over the course of the conversation, highlighting how user prompting can elicit better answers and underscoring the importance of multi-turn analysis beyond static benchmarks.
Even when the AI is not able to arrive at the correct answer, we find that users are sometimes still able to correct the mistake and select the right answer. We visualize the rates of these corrections in Figure B1.
@@ 5.1 Fine-Tuning a User Simulator
We then transform the real user-AI conversations from our study into training examples for supervised fine-tuning. Each conversation with $k$ user utterances yields $k+1$ training examples: one example in the Task 1 format where the gold standard response is the real user’s first utterance; k1k-1 examples in the Task 2 format where the gold standard response is each of the remaining utterances (providing the conversation up to that utterance); and one example in the Task 2 format with the full conversation and the gold standard response being “Answer: LETTER” corresponding to the user’s selected multiple choice option.
@@ Results.
As shown in Table 1, fine-tuning our simulator yields large gains, with a 22-26 point increase in correlation and a 21-26% decrease in MAE. As shown in Figures B3-B4, a primary failure mode of the simulator before fine-tuning is that it cannot replicate human mistakes and greatly overestimates user-AI performance, producing far more questions with accuracies of 100% than we see in the real user-AI distribution, while the fine-tuned simulator matches the real distribution more closely. We also find that fine-tuning improves the realism of the simulator’s generated utterances, with 11-16 point improvements in BLEU and ROUGE. The fine-tuned simulator also outperforms both AI-alone methods and IQA-EVAL across metrics.
@@ 8 Broader Impacts and Ethical Considerations
Our work is driven by broader impacts: we seek to make AI evaluation more realistic and human-centered, by investigating how evaluation conclusions change when we incorporate human interactions. With our user study, we show that evaluation conclusions change significantly from AI-alone to user-AI settings (for the same set of questions), and these results hold over different subject areas, AI models, AI-alone methods, and user-AI conditions. We hope that our work motivates AI researchers and practitioners to think more carefully about human-AI interactions when they evaluate AI systems, instead of only using AI-alone benchmarks.
@@ Acknowledgments
The authors thank Rich Ciapala for computing support, and Emma Pierson and Joseph Suh for helpful discussions.
@@ A.1 Task Details
We recruited workers on Prolific to participate in our study. All Prolific workers who were located in the US, fluent in English, and had not participated in one of our pilots were eligible for our study. We used Prolific’s standard sample, which distributed our study to available participants. Based on early pilots, we estimated that the task took around 25 minutes. We paid all participants \$5.00 upon completion of the entire task. We experimented with offering a small bonus per correct answer, which we discuss in Section A.2.
Participant demographics.
The demographic data we received from Prolific about the workers who participated in our study included their country of birth and residence, nationality, fluent languages, age, sex, ethnicity, and employment status. As a result of our eligibility criteria, all workers resided in the US and were fluent in English. We also found that the vast majority were born in the US (95% of the workers who participated in our full study), listed the US as their nationality (99%), and listed English as their primary language (99%).
Figure A1: Comparing the marginal distributions of sex, race, age, and employment status for Prolific workers (blue) vs US Census (orange). We only keep data in categories that appear in both datasets (e.g., we drop data values of DATA_EXPIRED or “Prefer not to say” in Prolific), which is why we sometimes have fewer than 650 Prolific workers (see $n$ in the legends). We then normalize the distributions over the kept categories to sum to 1.
Since the workers in our study were primarily from the US, we compare the other demographic variables to the demographic distribution of the US population. We acquire the joint distribution of sex, race, and age from the U.S. Census Bureau (2024), using their Monthly Postcensal Resident Population for July 2023 (the latest month before projections). For employment status, we use statistics from the U.S. Bureau of Labor Statistics (2025), using their estimates for February 2025.
In Figure A1, we compare each marginal distribution over the Prolific workers vs. the US population. The sex distributions are very well-matched: the proportions for "female" and "male" are (0.50, 0.50) on Prolific and (0.51, 0.49) in the US Census. The race distributions are also decently well-matched: using the categories provided to us by Prolific, the proportions for “White”, “Black”, “Asian”, “Mixed”, and “Other” are (0.72, 0.18, 0.04, 0.05, 0.02) on Prolific and (0.75, 0.14, 0.06, 0.03, 0.02) in the US Census. The employment status distribution is slightly different: the proportions for “Full-Time”, “Part-Time”, and “Unemployed (and job seeking)” are (0.68, 0.24, 0.07) on Prolific and (0.79, 0.17, 0.04) in US Labor Statistics, so individuals working full-time are slightly underrepresented while others are overrepresented. We see the largest divergence between distributions on age: individuals aged 25-34 are very overrepresented (0.34 vs. 0.17) while individuals aged 65 and above are very underrepresented (0.03 vs. 0.23). These trends are expected, given that younger individuals and those who are not employed full-time are likelier to work on crowdsourcing platforms Ipeirotis (2010); Posch et al. (2022); the younger lean of Prolific workers may also match the younger lean of generative AI users Bick et al. (2024). Even so, the non-representative nature of workers on crowdsourcing platforms should be taken into account as a limitation when interpreting our results.
Interface.
We implemented our app in React and used Azure CosmosDB as our database. We provide screenshots of all of the pages in our user study interface, including the Introduction Page (Figure A4), Phase 1 Tutorial (Figure A5), Confidence Page (Figure A6), User-Alone Page (Figure A7), Phase 2 Instructions (Figure A8), Phase 2 Tutorial (Figure A9), User-AI Page (Figure 2), and Feedback Page (Figure A10).
Dataset Model Comparison $\mathrm{Acc}_{1}$ $\mathrm{SE}_{1}$ $\mathrm{Acc}_{2}$ $\mathrm{SE}_{2}$ $z$-value $p$-value
Elementary Math GPT-4o AI letter zero shot vs. UserAI direct to ai 0.73 0.04 0.92 0.02 -3.92 <<0.01
Elementary Math GPT-4o AI letter zero shot vs. UserAI answer first 0.73 0.04 0.90 0.02 -3.43 <<0.01
Elementary Math GPT-4o AI letter few shot vs. UserAI direct to ai 0.74 0.04 0.92 0.02 -3.83 <<0.01
Elementary Math GPT-4o AI letter few shot vs. UserAI answer first 0.74 0.04 0.90 0.02 -3.34 <<0.01
Elementary Math GPT-4o AI free text vs. UserAI direct to ai 0.99 0.01 0.92 0.02 3.03 <<0.01
Elementary Math GPT-4o AI free text vs. UserAI answer first 0.99 0.01 0.90 0.02 4.04 <<0.01
Elementary Math GPT-4o User alone vs. UserAI direct to ai 0.78 0.03 0.92 0.02 -4.21 <<0.01
Elementary Math GPT-4o User alone vs. UserAI answer first 0.78 0.03 0.90 0.02 -3.52 <<0.01
High School Math GPT-4o AI letter zero shot vs. UserAI direct to ai 0.51 0.05 0.70 0.04 -3.20 <<0.01
High School Math GPT-4o AI letter zero shot vs. UserAI answer first 0.51 0.05 0.73 0.03 -3.92 <<0.01
High School Math GPT-4o AI letter few shot vs. UserAI direct to ai 0.49 0.04 0.70 0.04 -3.57 <<0.01
High School Math GPT-4o AI letter few shot vs. UserAI answer first 0.49 0.04 0.73 0.03 -4.33 <<0.01
High School Math GPT-4o AI free text vs. UserAI direct to ai 0.85 0.03 0.70 0.04 3.14 <<0.01
High School Math GPT-4o AI free text vs. UserAI answer first 0.85 0.03 0.73 0.03 2.73 <<0.01
High School Math GPT-4o User alone vs. UserAI direct to ai 0.41 0.03 0.70 0.04 -5.88 <<0.01
High School Math GPT-4o User alone vs. UserAI answer first 0.41 0.03 0.73 0.03 -7.03 <<0.01
College Math GPT-4o AI letter zero shot vs. UserAI direct to ai 0.45 0.07 0.52 0.08 -0.61 0.54
College Math GPT-4o AI letter zero shot vs. UserAI answer first 0.45 0.07 0.52 0.07 -0.72 0.47
College Math GPT-4o AI letter few shot vs. UserAI direct to ai 0.44 0.07 0.52 0.08 -0.72 0.47
College Math GPT-4o AI letter few shot vs. UserAI answer first 0.44 0.07 0.52 0.07 -0.85 0.40
College Math GPT-4o AI free text vs. UserAI direct to ai 0.73 0.06 0.52 0.08 2.23 0.03
College Math GPT-4o AI free text vs. UserAI answer first 0.73 0.06 0.52 0.07 2.40 0.02
College Math GPT-4o User alone vs. UserAI direct to ai 0.28 0.04 0.52 0.08 -2.67 <<0.01
College Math GPT-4o User alone vs. UserAI answer first 0.28 0.04 0.52 0.07 -3.10 <<0.01
Conceptual Physics GPT-4o AI letter zero shot vs. UserAI direct to ai 0.91 0.03 0.84 0.03 1.74 0.08
Conceptual Physics GPT-4o AI letter zero shot vs. UserAI answer first 0.91 0.03 0.84 0.03 1.70 0.09
Conceptual Physics GPT-4o AI letter few shot vs. UserAI direct to ai 0.96 0.02 0.84 0.03 3.22 <<0.01
Conceptual Physics GPT-4o AI letter few shot vs. UserAI answer first 0.96 0.02 0.84 0.03 3.22 <<0.01
Conceptual Physics GPT-4o AI free text vs. UserAI direct to ai 0.97 0.02 0.84 0.03 3.62 <<0.01
Conceptual Physics GPT-4o AI free text vs. UserAI answer first 0.97 0.02 0.84 0.03 3.63 <<0.01
Conceptual Physics GPT-4o User alone vs. UserAI direct to ai 0.55 0.03 0.84 0.03 -6.48 <<0.01
Conceptual Physics GPT-4o User alone vs. UserAI answer first 0.55 0.03 0.84 0.03 -6.69 <<0.01
Moral Scenarios GPT-4o AI letter zero shot vs. UserAI direct to ai 0.71 0.05 0.79 0.03 -1.47 0.14
Moral Scenarios GPT-4o AI letter zero shot vs. UserAI answer first 0.71 0.05 0.78 0.04 -1.13 0.26
Moral Scenarios GPT-4o AI letter few shot vs. UserAI direct to ai 0.80 0.04 0.79 0.03 0.27 0.79
Moral Scenarios GPT-4o AI letter few shot vs. UserAI answer first 0.80 0.04 0.78 0.04 0.49 0.63
Moral Scenarios GPT-4o AI free text vs. UserAI direct to ai 0.72 0.05 0.79 0.03 -1.26 0.21
Moral Scenarios GPT-4o AI free text vs. UserAI answer first 0.72 0.05 0.78 0.04 -0.93 0.35
Moral Scenarios GPT-4o User alone vs. UserAI direct to ai 0.73 0.03 0.79 0.03 -1.54 0.12
Moral Scenarios GPT-4o User alone vs. UserAI answer first 0.73 0.03 0.78 0.04 -1.05 0.29
Dataset Model Comparison $\mathrm{Acc}_{1}$ $\mathrm{SE}_{1}$ $\mathrm{Acc}_{2}$ $\mathrm{SE}_{2}$ $z$-value $p$-value
Elementary Math Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.45 0.04 0.86 0.03 -8.58 <<0.01
Elementary Math Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.45 0.04 0.90 0.02 -10.50 <<0.01
Elementary Math Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.43 0.03 0.86 0.03 -9.39 <<0.01
Elementary Math Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.43 0.03 0.90 0.02 -11.53 <<0.01
Elementary Math Llama-3.1-8b AI free text vs. UserAI direct to ai 0.88 0.03 0.86 0.03 0.56 0.58
Elementary Math Llama-3.1-8b AI free text vs. UserAI answer first 0.88 0.03 0.90 0.02 -0.65 0.51
Elementary Math Llama-3.1-8b User alone vs. UserAI direct to ai 0.81 0.03 0.86 0.03 -1.26 0.21
Elementary Math Llama-3.1-8b User alone vs. UserAI answer first 0.81 0.03 0.90 0.02 -2.70 <<0.01
High School Math Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.32 0.03 0.62 0.04 -6.14 <<0.01
High School Math Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.32 0.03 0.64 0.04 -6.89 <<0.01
High School Math Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.30 0.02 0.62 0.04 -7.09 <<0.01
High School Math Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.30 0.02 0.64 0.04 -7.98 <<0.01
High School Math Llama-3.1-8b AI free text vs. UserAI direct to ai 0.64 0.04 0.62 0.04 0.24 0.81
High School Math Llama-3.1-8b AI free text vs. UserAI answer first 0.64 0.04 0.64 0.04 -0.16 0.87
High School Math Llama-3.1-8b User alone vs. UserAI direct to ai 0.45 0.03 0.62 0.04 -3.37 <<0.01
High School Math Llama-3.1-8b User alone vs. UserAI answer first 0.45 0.03 0.64 0.04 -3.93 <<0.01
College Math Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.35 0.04 0.46 0.07 -1.37 0.17
College Math Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.35 0.04 0.48 0.07 -1.56 0.12
College Math Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.30 0.04 0.46 0.07 -1.97 0.05
College Math Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.30 0.04 0.48 0.07 -2.18 0.03
College Math Llama-3.1-8b AI free text vs. UserAI direct to ai 0.41 0.05 0.46 0.07 -0.57 0.57
College Math Llama-3.1-8b AI free text vs. UserAI answer first 0.41 0.05 0.48 0.07 -0.74 0.46
College Math Llama-3.1-8b User alone vs. UserAI direct to ai 0.40 0.04 0.46 0.07 -0.75 0.46
College Math Llama-3.1-8b User alone vs. UserAI answer first 0.40 0.04 0.48 0.07 -0.93 0.35
Conceptual Physics Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.53 0.05 0.67 0.04 -2.25 0.02
Conceptual Physics Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.53 0.05 0.73 0.04 -3.22 <<0.01
Conceptual Physics Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.57 0.04 0.67 0.04 -1.64 0.10
Conceptual Physics Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.57 0.04 0.73 0.04 -2.70 <<0.01
Conceptual Physics Llama-3.1-8b AI free text vs. UserAI direct to ai 0.62 0.04 0.67 0.04 -0.77 0.44
Conceptual Physics Llama-3.1-8b AI free text vs. UserAI answer first 0.62 0.04 0.73 0.04 -1.80 0.07
Conceptual Physics Llama-3.1-8b User alone vs. UserAI direct to ai 0.46 0.03 0.67 0.04 -3.91 <<0.01
Conceptual Physics Llama-3.1-8b User alone vs. UserAI answer first 0.46 0.03 0.73 0.04 -4.97 <<0.01
Moral Scenarios Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.40 0.03 0.72 0.04 -6.01 <<0.01
Moral Scenarios Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.40 0.03 0.74 0.04 -7.42 <<0.01
Moral Scenarios Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.31 0.03 0.72 0.04 -7.35 <<0.01
Moral Scenarios Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.31 0.03 0.74 0.04 -8.86 <<0.01
Moral Scenarios Llama-3.1-8b AI free text vs. UserAI direct to ai 0.49 0.03 0.72 0.04 -4.07 <<0.01
Moral Scenarios Llama-3.1-8b AI free text vs. UserAI answer first 0.49 0.03 0.74 0.04 -5.15 <<0.01
Moral Scenarios Llama-3.1-8b User alone vs. UserAI direct to ai 0.79 0.03 0.72 0.04 1.34 0.18
Moral Scenarios Llama-3.1-8b User alone vs. UserAI answer first 0.79 0.03 0.74 0.04 1.00 0.32
A.2 Pilots and Incentivization